You purchased a machine for $ 1.12 million three years ago and have been applying​ straight-line depreciation to zero for a​ seven-year life. Your tax rate is 40%. If you sell the machine today​ (after three years of​ depreciation) for $742,000​, what is your incremental cash flow from selling the​ machine?
Answer #1

COST OF MACHINE = 1120000

LIFE = 7 YEARS

DEPRECIATION PER YEAR = COST/LIFE = 1120000/ 7 = 160000

3 YEARS DEPRECIATION = 160000 X 3 = 480000

BOOK VALUE = COST - DEPRECIATION CHARGED FOR 3 YEARS

BOOK VALUE = 1120000 - 480000 = 640000

SALE VALUE OF MACHINE = 742000

GAIN ON SALE OF MACHINE = SALE VALUE - BOOK VALUE

GAIN ON SALE OF MACHINE = 742000 - 640000 = 102000

TAX ON GAIN ON SALE OF MACHINE = 102000 X 40% = 40800

INCREMENTAL CASH FLOW = SALE VALUE - TAX ON GAIN ON SALE OF MACHINE

INCREMENTAL CASH FLOW = 742000 - 40800 = 701200

Answer : 701200 (Thumbs up please)